James has 16 more game cards than Felix. If they have 48 game cards altogether, find the number of game cards James has.
Answer by ikleyn(52776)

.

James + Felix  = 48


J     + F      = 48


        F = J - 16,  according to the condition.


J     + (J-16) = 48


2J             = 48 + 16 = 64.


You continue from this point.


ANSWER.  James has 32 cards.